OpenStudy (anonymous):

what is the solution of the system of equations? -y + 3x = 6 y = -6x + 12

OpenStudy (texaschic101):

y = -6x + 12 now just sub -6x + 12 in for y in the other equation Let me get you started : -y + 3x = 6 -(-6x + 12) + 3x = 6 -- don't forget to distribute the negative through the parenthesis now just solve for x. Once x is found, sub that back into either equation and solve for y.